# AlphaOPT: Formulating Optimization Programs with Self-Improving LLM Experience Library

### v2

@@ Appendix B Datasets
We have collected the publicly available optimization problem datasets listed in the Table  4, which include both natural language problem descriptions and their optimal solutions. Among them, NLP4LP, NL4OPT, IndustryOR, and MAMO (ComplexLP) are used to construct our training and test datasets, while Optibench and LogiOR serve as out-of-distribution (OOD) evaluation sets. The gold-standard programs for the training datasets NLP4LP and IndustryOR are obtained from Yang et al. (2025a).
@@
Because our library-based framework derives knowledge feedback from correct solutions, it is relatively sensitive to data noise. Accordingly, we train and evaluate on clean splits that exclude instances labeled as erroneous, and the Size column in Table 4 indicates the size of each dataset after cleaning. Specifically, for NLP4LP, IndustryOR and LogiOR we use the cleaned versions provided by Yang et al. (2025a); for NL4OPT, MAMO (ComplexLP) and Optibench we use the cleaned releases from Astorga et al. (2025), obtained from their GitHub repository.
Table 4: The statistics of the optimization problem datasets
@@
Appendix C Proof of the Library Convergence
Recall the optimization problem in the library training phase
  1. 1.
    Therefore every accepted refinement strictly increases $F(\ell)$; otherwise the algorithm keeps $\ell_{k+1}=\ell_{k}$.
  2. 2.
    Sufficient exploration: Any improving neighbor $\tilde{\ell}\in R(\ell_{k})$ (i.e. one with strictly larger objective) will eventually be discovered and executed. Empirically, this is achieved through iterative prompt optimization with LLMs.
  3. 3.
    Boundedness: $r(t,\ell)$ and $\Omega(\ell)$ are bounded, hence $F(\ell)$ is bounded above and below.
Theorem 1.
Assume $\mathcal{T}_{\mathrm{train}}=\mathcal{T}_{\mathrm{test}}$. If the library space $\mathcal{L}$ is finite, then the algorithm terminates in finitely many steps at a library $\ell^{*}$ which is a local maximizer for the testing objective. Moreover, the algorithm cannot terminate at a saddle point.
Proof.
Every accepted merge or exploration step strictly increases $F(\ell)$, and otherwise the library remains unchanged. Since $F$ is bounded above, the sequence $\{F(\ell_{k})\}$ is monotone non-decreasing and bounded, hence convergent to some limit $F^{*}$. Furthermore, since $\mathcal{L}$ is finite, define
Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing. ∎
Theorem 2 (Infinite compact library case).
Proof.
Each accepted step strictly increases $F(\ell)$, so $\{F(\ell_{k})\}$ is monotone non-decreasing. Since $F$ is bounded above, $\{F(\ell_{k})\}$ converges to some $F^{*}$. By compactness of $\mathcal{L}$, there exists a convergent subsequence $\ell_{k_{j}}\to\ell^{\infty}$. Continuity of $F$ ensures $F(\ell_{k_{j}})\to F(\ell^{\infty})=F^{*}$. Suppose $\ell^{\infty}$ had a neighbor $\tilde{\ell}\in R(\ell^{\infty})$ with $F(\tilde{\ell})>F(\ell^{\infty})$. Then sufficient exploration would eventually yield $F(\ell_{k})>F^{*}$, which is a contradiction. Therefore, $\ell^{\infty}$ is a local maximizer. Saddle points are excluded. Since the training and testing distributions coincide, the training objective equals the testing objective, so any local optimality statement directly applies to testing. ∎
Theorems 1 and 2 together guarantee that, when the training and testing distributions coincide, the refinement algorithm converges to locally optimal solutions for the testing phase.
Appendix D Additional Results
D.1 Ablation Study
We assess the effectiveness of library insight retrieval/application and self-debug:
  • w/o self-debug: Remove the model’s self-debugging section.
  • w/o taxonomy: Remove the library taxonomy and directly match insights by checking all conditions.
  • w/o insight example: Use only the explanation as input, excluding exemplar cases
Table 5 shows that the full AlphaOPT achieves the best scores on all out-of-distribution datasets. Removing self-debug yields the largest drop on Logior (15.21%), indicating it plays an important role in the system. Dropping the library taxonomy reduces accuracy by 4.34% on Logior and by 1.09% on Optibench, suggesting that structured matching meaningfully improves retrieval. Excluding insight examples similarly lowers performance, showing that concrete, worked snippets aid application beyond textual explanations alone.
Table 5: The ablation results of AlphaOPT without: performance on benchmarks.
Dataset
Logior (92)
Optibench (403)
AlphaOPT (full) 51.08% 91.81%
w/o self-debug 35.87% 89.26%
w/o taxonomy 46.74% 90.72%
w/o insight example 45.65% 91.06%
D.2 Details of Library Insight Distribution
@@
Figure 8: Insight distributions across three different tracks, showing the contributions of source tasks in the four training datasets to the generation of library insights.
@@
In the Domain Modeling track, insights under Resource Allocation – Capacity/Resource Balance Equations account for the largest proportion, with contributions from all four datasets, among which NLP4LP and Industry OR contribute the most. Insights tagged with this label are widely applicable to optimization problems that require ensuring resource usage does not exceed available capacity, demonstrating how to establish constraints that maintain balance between resource consumption and availability.
@@
In the General Formulation track, Variable Definition – Explicit Bounds, Redundant Auxiliary Variable, and Continuous vs. Discrete Confusion are the most frequent, with diverse sources, indicating that these are common issues across multiple datasets. This reflects that foundational concepts in variable definition are the most error-prone in optimization modeling—particularly in balancing variable types, value ranges, and modeling simplifications—where LLMs tend to produce redundant formulations or type misuse due to overlooking structural or physical consistency.
@@
In the Code Implementation track, the number of insights is the smallest, with Solver & API Syntax – Nonlinear Constraint Handling and Data I/O & Validation – KeyError & Index Mismatch accounting for the highest proportions, mainly contributed by the MAMO (Complex LP) and NLP4LP datasets. These two types of insights reveal critical vulnerabilities in the implementation stage—solver syntax compatibility and data accessibility—representing the dual pillars required for bridging mathematical modeling and executable code.
@@
From the perspective of dataset contribution, Industry OR is dominant in Domain Modeling; MAMO (Complex LP) and Industry OR lead in General Formulation; and both MAMO (Complex LP) and NLP4LP contribute the most to Code Implementation. NL4OPT has relatively lower overall participation but focuses on formulation- and solver-related details. Considering dataset size, although the Industry OR and MAMO (Complex LP) datasets used for library learning are only about half the size of the other datasets, they still contribute a large number of insights, indicating that these datasets contain denser structural modeling challenges and more diverse error patterns, enabling the LLM to accumulate more experiential knowledge across multiple dimensions.
D.3 Examples of Library Refinement
We selected five representative refined insights from the library and explained how their applicability conditions were adjusted based on the associated tasks.
Figure 9: The examples of insight conditions before and after refinement, with green text indicating the original applicability conditions, and blue and red marking the newly added applicable and non-applicable conditions after refinement.
As shown in the Figure 9, Insight ID 10 targets single–scalar-capacity selection; the refinement broadens coverage by adding equivalent formulations such as minimizing a single resource (time/space), while explicitly excluding multi-capacity settings. Insight ID 43 captures flow/conservation and production balancing across nodes; refinement adds lexical anchors (e.g., capacity constraints, balancing production) and excludes statements dominated by fixed proportions, quality constraints, or predefined allocations, gating suppresses spurious retrievals where the structure is not true flow/conservation. Insight ID 100 supports adding tight explicit bounds derived from natural limits; refinement strengthens positive cues (e.g., maximum capacity, resource limits) and excludes cases that are purely discrete or already inherently bounded. Insight ID 133 focuses on max-flow between a single source and a single sink; refinement tightens the structural requirement and rules out source-to-many distribution tasks. Insight ID 228 covers translating at least/at most expressed as proportions into inequalities; refinement broadens to general percentage phrasing, clarifies the mapping, and excludes exact quantities, logical relations, or fixed ratios that are not proportion limits.
Across the five cases, four strategies proposed by LLM agent recur: (i) generalize with equivalent phrasings (e.g., minimize a single resource); (ii) lexical anchoring with positive keywords to raise recall where appropriate; (iii) explicit exclusion lists to reducing misalignment with tasks; (iv) structural qualifiers (single-scalar capacity; single-source–single-sink) to prevent misuse. A refinement is considered effective when the reduction in negative and unretrieved tasks outweighs the decrease in positive tasks. As shown in the Figure 10, these semantic refinements reduce both negative and unretrieved cases while preserving as many positive cases as possible.
Figure 10: Changes in the number of positive, negative, and unretrieved tasks associated with each insight before and after applicability condition refinement. A decrease in negative and unretrieved tasks indicates that, the insight no longer mismatches unrelated tasks (negative) and successfully matches previously applicable but missed tasks (unretrieved) after refinement.
D.4 Specification of Library Taxonomy Labels
@@
The following Table  6 lists all library taxonomy labels and their corresponding conditions, which specify the applicability criteria of each level-2 label and clarify its precise meaning. According to the library taxonomy generation mechanism, each label condition is created by the LLM when the label is first introduced, and subsequent generated insights under the same label inherit that initial condition.
Table 6: Full Specification of the Library Taxonomy
@@
Inventory Balance Equations Applies when the problem description involves materials or products that carry over between periods and must satisfy stock-flow balance.
@@
Appendix E Prompts For LLM Modules

### v3

@@ B.1. Datasets
We have collected the publicly available optimization problem datasets listed in Table  3, which include both natural language problem descriptions and their optimal solutions. Among them, NLP4LP, NL4OPT, IndustryOR, and MAMO (ComplexLP) are used to construct our training and test datasets, while Optibench and LogiOR serve as out-of-distribution (OOD) evaluation sets. The gold-standard programs for the training datasets NLP4LP and IndustryOR are obtained from Yang et al. (2025a).
@@
Because our library-based framework derives knowledge feedback from correct solutions, it is relatively sensitive to data noise. Accordingly, we train and evaluate on clean splits that exclude instances labeled as erroneous, and the Size column in Table 3 indicates the size of each dataset after cleaning. Specifically, for NLP4LP, IndustryOR, and LogiOR we use the cleaned versions provided by Yang et al. (2025a); for NL4OPT, MAMO (ComplexLP), and Optibench, we use the cleaned releases from Astorga et al. (2025), obtained from their GitHub repository.
Table 3. Statistics of the optimization problem datasets
@@
B.2. Micro- and Macro-Averaged Accuracy
Let $\mathcal{D}=\{D_{1},\dots,D_{K}\}$ be the set of evaluation datasets, with $|D_{k}|$ instances and accuracy $\mathrm{Acc}(D_{k})$. The micro-averaged accuracy is
$\mathrm{Acc}_{\text{micro}}=\frac{\sum_{k=1}^{K}|D_{k}|\,\mathrm{Acc}(D_{k})}{\sum_{k=1}^{K}|D_{k}|},$
which weights datasets by size. The macro-averaged accuracy is
$\mathrm{Acc}_{\text{macro}}=\frac{1}{K}\sum_{k=1}^{K}\mathrm{Acc}(D_{k}),$
which weights all datasets equally.
B.3. Insights Learned Through Baseline Experience Learning Methods
To better understand the limitations of prior experience-learning methods in optimization formulation, we examine all insights extracted by ExpeL when adapted to operations research tasks. The learned insights are dominated by generic, high-level recommendations, listed below:
  • Ensure that decision variables are defined with the correct type (integer, binary, or continuous) based on the problem requirements, especially for inherently discrete quantities.
  • Double-check that all constraints are correctly implemented and aligned with the problem conditions, including the correct mathematical representation and direction of inequalities.
  • Verify that the objective function accurately reflects the intended optimization goal.
  • Always validate the model output against expected results to ensure feasibility and correctness.
  • Regularly review the problem statement to ensure correct interpretation of variable types and constraints.
  • Ensure that decision variables have appropriate bounds to reflect realistic limits and prevent infeasible or unbounded solutions.
  • Incorporate logical constraints to model conditional relationships between decision variables.
  • Test the model under different scenarios to assess robustness and adaptability.
  • Re-validate variable types when transitioning between continuous and discrete representations.
  • Check the model structure for potential infeasibilities or unboundedness before optimization.
  • Consider edge cases to improve robustness across problem conditions.
  • Ensure that the objective and constraints scale appropriately with problem size or parameter changes.
  • Examine the impact of variable bounds on feasibility and correctness.
These insights emphasize general modeling hygiene rather than concrete modeling or coding practices. They do not encode explicit applicability conditions, domain-specific structural patterns (e.g., flow conservation, inventory balance, fixed-charge activation), or solver-level implementation details. Consequently, while such insights are broadly reusable, they provide limited actionable guidance for correcting specific formulation or implementation errors in complex optimization problems.
B.4. Experimental Cost
All services used in this work rely on third-party API calls to OpenAI’s GPT-4o models. We report the token consumption incurred during both the training and evaluation phases of AlphaOPT. Here Steps denotes the number of execution rounds required by each module or phase. In practice, the iterative Library Diagnosis–Refinement circle in AlphaOPT typically converges within five rounds during training.
As shown in the table, the Library Refinement phase consumes more prompt tokens, since refining each insight requires incorporating all of its associated tasks as contextual input. During the testing phase, the token cost is only related to solution generation and insights resue, resulting in a substantially lower token consumption compared to the training stage.
Overall, under well-designed prompt engineering and constrained execution steps, the average cost per instance remains around \$ 0.1, and the total experimental cost is kept at a moderate and acceptable level. Moreover, owing to the reusability of the library, the relatively high computational cost is incurred only once during the construction of experience library; therefore, the overall system exhibits strong economic scalability in large-scale testing or long-term usage scenarios.
Table 4. Computation cost and runtime statistics across training and testing stages. (to be updated
Stage Phase Data Size Prompt Tokens Completion Tokens Cost per instance Runtime Steps
Training Library Learning 452 $\sim$7.7M $\sim$1.3M $\sim$\$0.07 $\sim$40.5 min 1
Training Library Diagnosis 452 $\sim$11.5M $\sim$2.3M $\sim$\$0.11 $\sim$48.6 min 3-5
Training Library Refinement 452 $\sim$32.9M $\sim$1.9M $\sim$\$0.22 $\sim$33.7 min 3-5
Testing 691 $\sim$13.8M $\sim$1.5M $\sim$\$0.07 $\sim$25.5 min 3
B.5. Insight Distribution
@@
Figure 7. Insight distributions across three different tracks, showing the contributions of source tasks in the four training datasets to the generation of library insights.
The figure shows how insights in the experience library are distributed across the three main tracks and how different source datasets contribute to each track. Panel (a) focuses on the Domain Modeling track. Each horizontal bar corresponds to a specific domain-level modeling concept, such as capacity and resource balance equations, flow conservation, max-flow objectives, or inventory balance. The bars are segmented by dataset, indicating how many insights of each type originate from NLP4LP, Industry OR, MAMO Complex, and NL4OPT tasks. This panel highlights that domain modeling insights are concentrated around a small number of common structures, especially resource allocation and network flow patterns, with contributions coming from multiple datasets. Panel (b) shows the distribution for the General Formulation track. The listed concepts include variable definition issues, constraint formulation errors, unit scaling problems, and objective specification confusions. Compared to domain modeling, these insights are more evenly spread across formulation-level pitfalls, such as continuous versus discrete variables, redundant auxiliary variables, and Big-M related issues. Contributions again vary across datasets, reflecting differences in how formulation errors arise in different task sources. Panel (c) presents the Code Implementation track. The insights here relate to solver and API usage, constraint syntax, variable bounds, and data input or validation errors. The distribution shows fewer categories but clear concentration on solver syntax and implementation mistakes, with notable contributions from datasets that emphasize executable optimization code. Overall, the figure illustrates that different datasets contribute complementary types of insights and that the learned library captures a diverse range of issues spanning domain modeling, mathematical formulation, and code-level implementation.
@@
In the Domain Modeling track, insights under Resource Allocation – Capacity/Resource Balance Equations account for the largest proportion, with contributions from all four datasets, among which NLP4LP and Industry OR contribute the most. Insights tagged with this label are widely applicable to optimization problems that require ensuring that resource usage does not exceed available capacity. This demonstrates how to establish constraints that maintain balance between resource consumption and availability.
@@
In the General Formulation track, Variable Definition – Explicit Bounds, Redundant Auxiliary Variable, and Continuous vs. Discrete Confusion are the most frequent, with diverse sources, which indicates that these are common issues across multiple datasets. This reflects the fact that foundational concepts in variable definition are the most error-prone in optimization modeling—particularly in balancing variable types, value ranges, and modeling simplifications—while LLMs tend to produce redundant formulations or type misuse due to overlooking structural or physical consistency.
@@
In the Code Implementation track, the number of insights is the smallest, with Solver & API Syntax – Nonlinear Constraint Handling and Data I/O & Validation – KeyError & Index Mismatch accounting for the highest proportions, which are mainly contributed by the MAMO (Complex LP) and NLP4LP datasets. These two types of insights reveal critical vulnerabilities in the implementation stage—solver syntax compatibility and data accessibility—and represent the dual pillars required for bridging mathematical modeling and executable code.
@@
From the perspective of dataset contribution, Industry OR is dominant in Domain Modeling; MAMO (Complex LP) and Industry OR lead in General Formulation; and both MAMO (Complex LP) and NLP4LP contribute the most to Code Implementation. NL4OPT has relatively lower overall participation but focuses on formulation- and solver-related details. Considering dataset size, although the Industry OR and MAMO (Complex LP) datasets used for library learning are only about half the size of the other datasets, they still contribute a large number of insights, which indicates that these datasets contain denser structural modeling challenges and more diverse error patterns; this enables the LLM to accumulate more experiential knowledge across multiple dimensions.
B.6. Success and Failure Case Study
To demonstrate the effectiveness of the retrieved library insights, we perform case studies across all evaluation datasets (including the test splits of NLP4OPT, NLP4LP, IndustryOR, and MAMO-complex, as well as the out-of-distribution sets LogiOR and Optibench) in which we compare the performance of solving optimization tasks with and without library retrieval enabled, while excluding tasks that succeed in both settings from the evaluation. As shown in Figure 8, approximately half of the retrieved insights successfully contributed to solving new optimization tasks, with only a small portion introducing new errors (Failure). This demonstrates the effectiveness of our method for automating knowledge transfer for optimization modeling. Moreover, insights under different level-1 taxonomy labels exhibit distinct performance patterns across modeling dimensions.
(a) Overall effectiveness of retrieved library insights
(b) Effectiveness distribution across Level-1 insight categories
Figure 8. We evaluate the effectiveness of retrieved insights by classifying their outcomes as success, failure, or invalid. An insight is considered successful if it matches a new task and prevents an error that would otherwise occur, failed if it mismatches a task and introduces a new error, and invalid if it matches correctly but fails to help the LLM fix the original mistake. (a) The overall distribution of insight outcomes. (b) The proportions of successful, failed, and invalid insights under level-1 taxonomy labels.
The figure evaluates how effective the retrieved library insights are when applied to new tasks. Panel (a) shows the overall outcome distribution for all retrieved insights. Each insight is classified as a success, failure, or invalid. Successful cases are those where the insight matches the task and prevents an error that would otherwise occur. Failed cases are those where the insight mismatches the task and introduces a new error. Invalid cases are those where the insight is applicable but does not help correct the original mistake. The pie chart reports both counts and proportions, showing that nearly half of the retrieved insights lead to successful corrections, while the rest are split between invalid and failed outcomes. Panel (b) breaks down these outcomes by level-1 insight categories. Each horizontal bar corresponds to a category such as domain modeling, general formulation, or code implementation, and is further divided into success, failure, and invalid portions. This view highlights differences in effectiveness across categories. For example, variable definition and solver or API syntax insights account for a large number of successful cases, while some domain modeling categories show higher proportions of invalid outcomes. Overall, the figure shows that retrieved insights are often helpful, with some variation between categories.
1) Insights with high success rates
These insights are concentrated in Code Implementation – Solver & API Syntax and General Formulation – Variable Definition. These insights typically target code implementation or fundamental modeling errors with clear structures, which enable the LLM to follow their guidance stably and correctly. For instance, insights tagged with the Strict Inequality label under Solver & API Syntax highlight the fact that solvers (e.g., Gurobi) do not support strict inequalities and should instead be reformulated as non-strict forms ($\leq-1$ or $\geq+1$); insights tagged with the Explicit Bounds label under Variable Definition emphasize that decision variables should be assigned explicit upper and lower bounds to ensure feasibility and improve solver efficiency.
2) Insights with high failure rates
These insights are mainly found in Domain Modeling – Facility Location. They often involve structural constraints and logical triggers that are easily misinterpreted or overgeneralized by the LLM. For example, insights tagged with the Fixed Charge (Big-M Linking) label exhibit a failure rate of 70%. Although the principle of using Big-M constraints to model on/off logic is correct, its blind application in problems without conditional activation can lead to redundant or overlapping constraints and unnecessary feasible-region reduction, which ultimately causes solver failure. Also, while insights under Variable Definition generally achieve high success rates, some insights tagged with Explicit Bounds label sometimes are overly rigid, which leads the LLM to impose unnecessary upper bounds and thereby restricts the feasible space and produces suboptimal solutions.
3) Insights with high invalid rates
These insights are mainly under Domain Modeling – Resource Allocation and Domain Modeling – Product Planning. Although the LLM successfully retrieves the correct insights and identifies the corresponding problem types, it often fails to translate them into executable formulas or solver-level implementations. For instance, under Solver & API Syntax, Nonlinear Constraint Handling advises linearization or the introduction of auxiliary variables for nonlinear objectives or constraints (e.g., ratios or divisions), yet the LLM frequently fails to fully execute these transformations (by neglecting auxiliary variables or mis-rewriting proportional constraints), which results in insights being recognized but not operationalized.
Also, certain tasks remained unsolved regardless of whether library retrieval was enabled, because their failure stemmed from factors beyond the current learned library’s knowledge scope. In the out-of-distribution dataset LogiOR, tasks involved multi-level spatiotemporal logic and interacting constraints (e.g., capacity, timing, and flow balance) in problems such as routing, scheduling, and network flow. These challenges extend beyond the scope of the existing library, which primarily focuses on static, linear formulations. Although related taxonomy labels such as Resource Allocation and Nonlinear Constraint Handling exist, their granularity and depth are insufficient for modeling such complex logic. This demonstrates that the current system, while semantically generalizable, still lacks robust cross-structural transfer and context adaptation capabilities.
B.7. Specification of Library Taxonomy Labels
@@
The following Table  5 lists all library taxonomy labels and their corresponding conditions, which specify the applicability criteria of each level-2 label and clarify its precise meaning. According to the library taxonomy generation mechanism, each label condition is created by the LLM when the label is first introduced, and subsequent insights generated under the same label inherit that initial condition.
Table 5. Full Specification of the Library Taxonomy
@@
Inventory Balance Equations Applies when the problem description involves materials or products that carry over between periods and must satisfy the stock-flow balance.
@@
Appendix C Prompts For LLM Modules